OpenStudy (anonymous):

4) If h(x) =1/2 x^4 - 3x^3, find h(4). A) -200 B) -64 C) -28 D) -16

OpenStudy (campbell_st):

h(4) = 1/2 4^4 - 3 4^3 = 128 - 192 =- 64 or -4^3

OpenStudy (anonymous):

Thank you